OpenStudy (joshawad):

.03y+1=4.3 -.03 -.03 Y+1=4.27 Y=1-4.27 Y=3.27 Is this correct?

OpenStudy (joshawad):

@texaschic101 help again please

OpenStudy (texaschic101):

no... .03y + 1 = 4.3 you want to get the y term on one side by itself and everything else on the other side .03y + 1 = 4.3 -- subtract 1 from both sides .03y + 1 - 1 = 4.3 - 1 -- combine like terms .03y = 3.3 -- now you have to divide by .03 to get y by itself .03y/.03 = 3.3/.03 y = 110

OpenStudy (joshawad):

oh yeah i forgot that i have to divide to seperate numbers from variables

OpenStudy (joshawad):

how did u get 110 i got 1.1

OpenStudy (joshawad):

@texaschic101

OpenStudy (texaschic101):

3.3/.03 = 110

OpenStudy (joshawad):

oh sorry i div 3.3/.3 not .03 thanks again!!!

OpenStudy (texaschic101):

no problem...you got to watch those decimals, they can mess you up if you forget them or put them in the wrong spot...lol